# Diagrammatic expansion for the mutual-information rate in the realm of limited statistics

### v2

@@ II Background
with $S^{\mathrm{out}}\left(\Delta t\right)$ the output (or marginal) entropy and the input (or conditional) entropy $S_{\xi}^{\mathrm{in}}\left(\Delta t\right)$ conditional on the stimulus. The choice of $\Delta t$ should depend on the system dynamics, and in particular on the relevant time scale of the stimulus. Yet, the dynamics of biological systems as the retina extends over multiple time scales, and cannot be captured by a single time bin. In order to understand the consequences of this choice, we consider the toy example of a neuron firing according to an inhomogeneous Poisson process. Mimicking the effect of a dynamical stimulus, the neuron’s rate randomly switches between a low and high state with an average frequency of $27\,\mathrm{Hz}$, cf. 1a. This yields an exponential decay of the autocorrelation of the neuron, both for the mean activity over repetitions (peristimulus time histogram, PSTH) and for its spike count (respectively, blue dashed and black line in 1b). Using equation (1) to compute the MIR for this process leads to an estimate monotonously decaying with the time-bin size (1c), showing that the choice of the bin duration strongly affects the MIR estimation. For large $\Delta t$ the stimulus dynamics is averaged out, and the MIR vanish. Upon decreasing $\Delta t$ this effect reduces, with the estimate for the mutual-information rate monotonously attaining a limiting value for $\Delta t\rightarrow 0$. However, this is not a generic behavior, as we figure out by investigating the same system, but after adding a refractory period to the model neuron, i.e. it stays silent for $10\,\mathrm{ms}$ after every spike, cf. 1d. The other parameters are unchanged. This leads to lower firing rates, more regular spike trains and spike-count autocorrelations which are negative and large for small times (1e). Refractory periods decrease the neurons’ variability, therefore increasing their capacity of transmitting information at fixed firing rate [25]. In order for the effect of the refractory period to become noticeable in the $\mathrm{MIR}$, however, $\Delta t$ has to be large enough (for a more detailed explanation see A.2 in the appendix). Consistently, the naïve MIR is not monotonic anymore and shows a maximum at around $15\,\mathrm{ms}$ (1f). While in the limit of $\Delta t\rightarrow 0$ we avoid stimulus averaging, we will neglect the positive effect of refractoriness, leading to an underestimation of the MIR. So, already in the case with only two time scales, there is no good choice for the length of $\Delta t$ fully accounting for the system dynamics.
@@
Figure 1: (a) A neuron emitting spikes according to a rate switching randomly (Poisson process) between two levels, whereas the spiking conditional on the rate is either Poissonian itself (also panels b and c) or (d) features a refractory period (also panels e and f), a dot indicates a spike. (b,e) Autocorrelations of the mean activity indicated by the dotted blue lines, autocorrelations of the spikes by the solid black lines. (c,f) Mutual-information rates computed lumping activity into one time bin, as in equation (1), as a function of the time-bin size $\Delta t$. Parameters: Correlation time of switching of the rate: $T_{\mathrm{switch}}=100\,\mathrm{ms}$, firing rate $f_{\mathrm{low}}=5\mathrm{Hz}$, $f_{\mathrm{high}}=50\mathrm{Hz}$, (absolute) refractory period $t_{\mathrm{ref.}}=10\,\mathrm{ms}$, simulation time $T_{\mathrm{total}}=3\cdot 10^{4}\,\mathrm{ms}$, $N_{\mathrm{rep.}}=10^{4}$.
@@
Figure 2: (a) Sketch of a recording of spiking neurons responding to a repeated stimulus over a time span $\Delta t$ with the activity discretised into bins of length $\mathrm{dt}$. (b) estimates of the $\mathrm{MIR}$ as equation (3), computed by the (mixed) Moba-MIRA method, introduced in III, for varying recording time $\Delta t$, along with their naive estimates, equation (1), as shown in 1, panel c. Time-bin size $\mathrm{dt}=20\mathrm{ms}$. (c) As panel b, but with refractory period, as in 1f. Other parameters as in 1.
The required entropies for the $\mathrm{MIR}$s as shown in 2 are those of probability distributions of stochastic paths across multiple time bins. They are high-dimensional objects because $\Delta t$ has to be large enough to cover the correlation time of the stimulus (and the response) and $\mathrm{dt}$ has to be small enough to capture the dynamics of the neural activity. For a limited amount of data, as is typical for experiments, these entropies quickly become difficult or impossible to compute reliably. The direct method to determine them, namely counting the occurrence of patterns (histogram method [7]) can therefore only work in simple setups, in particular for single neurons [6, 12, 13, 14, 15]. Even then, the number of time bins is limited (up to eight in the cited examples) because the number of possible words grows exponentially in this quantity. This issue applies in particular to the input entropy (conditional on the stimulus, cf. equation(2)), because it quantifies the variability given a stimulus. It can therefore only be measured by repeating the same stimulus several times and collecting statistics over these repetitions. Here, therefore, number of samples equals that of the repetitions of the same stimulus, so typically at most about 100 in real data. In contrast, the output entropy marginalizing over the stimulus is computed over all times and repetitions, therefore the underlying probability distribution is much better sampled.
@@ III.1 Maximum-entropy modeling
where $S_{0}$ is the entropy of all single bins summed up, neglecting correlations, $c$ is the covariance matrix between bins (the autocovariance of the neuron under scrutiny across time) and $V$ are the respective variances, written as the entries of a diagonal matrix. This is what we call the moment-based mutual-information-rate approximation (Moba-MIRA). More precisely, we will use two versions of it: in one, we will use equation (4) only for the input entropy, while estimating the output entropy by the histogram method. We will call this variant the mixed Moba-MIRA, whereas we christen full Moba-MIRA the variant for which we use equation (4) for both types of entropies.
The approximation of equation (4) makes sense intuitively: it is nearly the Gaussian approximation, but with the important difference that we exactly take into account the entropies of the activitities (not necessarily binary) in single time bins. In the following we derive equation (4) in more grounded way using results from statistical physics and a diagrammatic expansion and explain how to add additional higher-order corrections.
@@ III.1.1 Derivation of the Moba-MIRA approximation
One possibility to use equation (5) would now be to fit the parameters $J_{tt^{\prime}}$ and in $H_{t}$ and then sample from $P$ to compute the entropy. However, we can get around this step by leveraging recently developed techniques from statistical field theory [18, 19, 20], explained in more detail in app. A.1. They allow us to derive the approximation equation (4) as a resummation of a class of diagrams in the diagrammatic small-correlation expansion around the case of uncorrelated time bins
@@
which, using Feynman-diagrammatic rules  [^2] In statistical physics one sometimes also refers to them as Mayer diagrams [50, 51]. The work of Mayer & Goeppert-Mayer is actually a few years older than that of Feynman. indeed yields equation (4). Note that the elements of the diagrams represent combinations of the measured cumulants - we do not have to know $J_{tt^{\prime}}$ or $H_{t}$ explicitly. We will sketch the basics of diagrammatics in appendix A.1 and refer to [18, 19, 20] for a more detailed description. Considering corrections to equation (LABEL:eq:Approx_SingleBinEnt_Loops) by taking into account more diagrams can be beneficial in some cases, but we have found the resummed-loop approximation to be the most robust.
@@ III.2 Testing the approximation on artificial data
How does the dependence of $\mathrm{MIR}$ estimates on $\mathrm{dt}$ and $\Delta t$ in the regime of limited data looks like for other methods? In 4a, we demonstrate that assuming the data to be simply Gaussian, which amounts to setting $S=\ln\left(\det\left(c\right)\right)+\mathrm{const.}$ leads to a clear overestimation of the $\mathrm{MIR}$ (yellow bar). Also, just neglecting correlations between time bins is not feasible, as the green bar shows, indicating that actually about half of the information is captured by the interaction between bins. Characterizing this relation using the histogram method, assuming a quadratic dependence on $1/N_{R}$ and extrapolating to $1/N_{R}=0$, as suggested by Strong et al. [7], yields much better results (orange bar), but leads still to an overestimate. However, assuming a probability distribution of the form equation (5) for the probability distribution conditional on the stimulus and applying the approximation equation (4) (mixed Moba-MIRA) yields an excellent fit, provided that we remove the bias by subtracting the estimates obtained from shuffled data, see A.5 for details. Imposing the form of equation (5) for the output entropy as well (full Moba-MIRA) yields a slightly worse fit, however, still performs better than the method from [7].
Why does our approach, equation (4), work? In general pairwise probability distributions, equation (5), might not be appropriate because the data could be also shaped by higher-order interactions in addition. However, for the input entropy, the most important part of the covariance is due to the refractory period of the neurons, as visible in the auto-correlation of the spikes, cf. A1. In other words, given the stimulus, the activity in two time bins is correlated mostly because a spike in one time bin suppresses a spike in another one  [^3] Note also that this interaction is not directed in time - whenever there is a spike at some point in time, one knows that there could not have been a spike both before and after this time. This effective suppression of spikes in neighbored time bins is an intrinsically two-point like interaction - at least as long as the time bin is not considerably shorter than the refractory period. It therefore does not come as a surprise that this approach works well for the input entropy. For the output entropy, however, we cannot make a similar argument and indeed, the pairwise approach clearly works worse in this case, in particular for smaller time bins.
@@ IV Results: Application of Moba-MIRA on retinal stimulus response
The results we have presented so far were obtained with the “full” version of Moba-MIRA, that is, we have used the approximation equation (4) for the input and the output entropies. However, in the analysis of the artificial data, we have seen that mixed Moba-MIRA yields results visibly closer to the ground truth, cf. 4a. Also, we have found single ON-cells for which full Moba-MIRA seems to deliver unreliable results, see A.5. Luckily, the estimates for the $\mathrm{MIR}$ of our real data typically already converge at values for $\Delta t$ as low as $80\mathrm{ms}$, see A2. In this case $k=8$ so that it is feasible to compute the estimate for $\mathrm{MIR}$ also in the mixed-Moba-MIRA version. This allows us to compare our results shown in 5 and 6, obtained by full Moba-MIRA, by their counterparts obtained by mixed Moba-MIRA, which we show in III.2. We observe that there are no qualitative differences between these two approaches.
@@ V Discussion
Also, the pairwise model equation (5) provides even more potential to reduce numerical costs: By using the alternative definition
@@
for the mutual-information rate, one can reduce the numerical cost from being quadratic in $k$ to linear. Indeed, with the definition equation (3), one has to compute the correlations of $\frac{k\left(k-1\right)}{2}$ pairs of bins, assembled from the $k$ time bins in the time window $\Delta t$. With the definition equation (7), however, one only has to compute the correlation of the activity in the most advanced time bin, corresponding to the interval $\left(\Delta t,\Delta t+\mathrm{dt}\right]$, with the $k$ bins inside $\left[0,\Delta t\right]$, therefore only $k$ correlations. As noted before [11], both definitions for the MIR are equivalent when the correlations between time bins decay to 0. In practice, the achievable $\Delta t$ might not always be large enough to make both expressions agree, but even then, equation (7) is an appealing alternative because it is closer to a differential definition of a rate or flux.
@@
In this work we applied Moba-MIRA to estimate mutual-information rates of individual neurons. Our method can however be extended to account for populations by modeling the correlation between different neurons at different times. Even if undersampling might be an issue there, we expect Moba-MIRA to be very useful, as methods based on histogram approaches would require an even larger amount of data, often beyond existing experimental datasets. With Moba-MIRA, however, only the estimation of correlations matrices is required, which reduces the necessary dataset size. We thus expect that reliable estimates can be given at least for pairs of neurons, analyzed with a temporal resolution comparable to that employed in this study. Currently, extending Moba-MIRA in that direction can be hindered by the lack of ground truth estimation for large populations, and because of this we leave it for future developments.
@@ Appendix A Technical details and additional analysis
where by $\left\langle\dots\right\rangle_{P}$ we denote the average over the empirical distribution. The function $H_{i}$ is arbitrary in our technical framework - one might choose, e.g., $H_{i}\left(n\right)=-\ln\left(n!\right)-e^{h_{i}}$ in case one would like to expand around a Poissonian theory. We, instead, implement a maximum-entropy approach choosing it as power series according to equation (6). With this choice, we force our statistical model to reproduce all empirically measured cumulants for single bins. This is ensured by choosing the interaction matrix $J$ and the (infinitely many) parameters of $\boldsymbol{H}$, $\lambda_{1},\lambda_{2},\dots$ accordingly. Because $\ln\left({\cal Z}\right)$ is the cumulant-generating function, we can express this condition as

### v3

@@ II Background
with $S^{\mathrm{out}}\left(\Delta t\right)$ the output (or marginal) entropy and the input (or conditional) entropy $S_{\xi}^{\mathrm{in}}\left(\Delta t\right)$ conditional on the stimulus. The choice of $\Delta t$ should depend on the system dynamics, and in particular on the relevant time scale of the stimulus. Yet, the dynamics of biological systems as the retina extends over multiple time scales, and cannot be captured by a single time bin. In order to understand the consequences of this choice, we consider the toy example of a neuron firing according to an inhomogeneous Poisson process. Mimicking the effect of a dynamical stimulus, the neuron’s rate randomly switches between a low and high state with an average frequency of $27\,\mathrm{Hz}$, cf. 1a. This yields an exponential decay of the autocorrelation of the neuron, both for the mean activity over repetitions (peristimulus time histogram, PSTH) and for its spike count (respectively, blue dashed and black line in 1b). Using (LABEL:Def_MIR_naiv) to compute the MIR for this process leads to an estimate monotonously decaying with the time-bin size (1c), showing that the choice of the bin duration strongly affects the MIR estimation. For large $\Delta t$ the stimulus dynamics is averaged out, and the MIR vanish. Upon decreasing $\Delta t$ this effect reduces, with the estimate for the mutual-information rate monotonously attaining a limiting value for $\Delta t\rightarrow 0$. However, this is not a generic behavior, as we figure out by investigating the same system, but after adding a refractory period to the model neuron, i.e. it stays silent for $10\,\mathrm{ms}$ after every spike, cf. 1d. The other parameters are unchanged. This leads to lower firing rates, more regular spike trains and spike-count autocorrelations which are negative and large for small times (1e). Refractory periods decrease the neurons’ variability, therefore increasing their capacity of transmitting information at fixed firing rate [25]. In order for the effect of the refractory period to become noticeable in the $\mathrm{MIR}$, however, $\Delta t$ has to be large enough (for a more detailed explanation see A.2 in the appendix). Consistently, the naïve MIR is not monotonic anymore and shows a maximum at around $15\,\mathrm{ms}$ (1f). While in the limit of $\Delta t\rightarrow 0$ we avoid stimulus averaging, we will neglect the positive effect of refractoriness, leading to an underestimation of the MIR. So, already in the case with only two time scales, there is no good choice for the length of $\Delta t$ fully accounting for the system dynamics.
@@
Figure 1: (a) A neuron emitting spikes according to a rate switching randomly (Poisson process) between two levels, whereas the spiking conditional on the rate is either Poissonian itself (also panels b and c) or (d) features a refractory period (also panels e and f), a dot indicates a spike. (b,e) Autocorrelations of the mean activity indicated by the dotted blue lines, autocorrelations of the spikes by the solid black lines. (c,f) Mutual-information rates computed lumping activity into one time bin, as in (LABEL:Def_MIR_naiv), as a function of the time-bin size $\Delta t$. Parameters: Correlation time of switching of the rate: $T_{\mathrm{switch}}=100\,\mathrm{ms}$, firing rate $f_{\mathrm{low}}=5\mathrm{Hz}$, $f_{\mathrm{high}}=50\mathrm{Hz}$, (absolute) refractory period $t_{\mathrm{ref.}}=10\,\mathrm{ms}$, simulation time $T_{\mathrm{total}}=3\cdot 10^{4}\,\mathrm{ms}$, $N_{\mathrm{rep.}}=10^{4}$.
@@
Figure 2: (a) Sketch of a recording of spiking neurons responding to a repeated stimulus over a time span $\Delta t$ with the activity discretised into bins of length $\mathrm{dt}$. (b) estimates of the $\mathrm{MIR}$ as (LABEL:MIR_def_proper), computed by the (mixed) Moba-MIRA method, introduced in III, for varying recording time $\Delta t$, along with their naive estimates, (LABEL:Def_MIR_naiv), as shown in 1, panel c. Time-bin size $\mathrm{dt}=20\mathrm{ms}$. (c) As panel b, but with refractory period, as in 1f. Other parameters as in 1.
The required entropies for the $\mathrm{MIR}$s as shown in 2 are those of probability distributions of stochastic paths across multiple time bins. They are high-dimensional objects because $\Delta t$ has to be large enough to cover the correlation time of the stimulus (and the response) and $\mathrm{dt}$ has to be small enough to capture the dynamics of the neural activity. For a limited amount of data, as is typical for experiments, these entropies quickly become difficult or impossible to compute reliably. The direct method to determine them, namely counting the occurrence of patterns (histogram method [7]) can therefore only work in simple setups, in particular for single neurons [6, 12, 13, 14, 15]. Even then, the number of time bins is limited (up to eight in the cited examples) because the number of possible words grows exponentially in this quantity. This issue applies in particular to the input entropy (conditional on the stimulus, cf. (LABEL:Def_MutualInfo)), because it quantifies the variability given a stimulus. It can therefore only be measured by repeating the same stimulus several times and collecting statistics over these repetitions. Here, therefore, number of samples equals that of the repetitions of the same stimulus, so typically at most about 100 in real data. In contrast, the output entropy marginalizing over the stimulus is computed over all times and repetitions, therefore the underlying probability distribution is much better sampled.
@@ III.1 Maximum-entropy modeling
where $S_{0}$ is the entropy of all single bins summed up, neglecting correlations, $c$ is the covariance matrix between bins (the autocovariance of the neuron under scrutiny across time) and $V$ are the respective variances, written as the entries of a diagonal matrix. This is what we call the moment-based mutual-information-rate approximation (Moba-MIRA). More precisely, we will use two versions of it: in one, we will use (LABEL:Entropy_single_plus_loops) only for the input entropy, while estimating the output entropy by the histogram method. We will call this variant the mixed Moba-MIRA, whereas we christen full Moba-MIRA the variant for which we use (LABEL:Entropy_single_plus_loops) for both types of entropies.
The approximation of (LABEL:Entropy_single_plus_loops) makes sense intuitively: it is nearly the Gaussian approximation, but with the important difference that we exactly take into account the entropies of the activitities (not necessarily binary) in single time bins. In the following we derive (LABEL:Entropy_single_plus_loops) in more grounded way using results from statistical physics and a diagrammatic expansion and explain how to add additional higher-order corrections.
@@ III.1.1 Derivation of the Moba-MIRA approximation
One possibility to use (LABEL:Def_generalized_Ising) would now be to fit the parameters $J_{tt^{\prime}}$ and in $H_{t}$ and then sample from $P$ to compute the entropy. However, we can get around this step by leveraging recently developed techniques from statistical field theory [18, 19, 20], explained in more detail in app. A.1. They allow us to derive the approximation (LABEL:Entropy_single_plus_loops) as a resummation of a class of diagrams in the diagrammatic small-correlation expansion around the case of uncorrelated time bins
@@
which, using Feynman-diagrammatic rules  [^2] In statistical physics one sometimes also refers to them as Mayer diagrams [53, 54]. The work of Mayer & Goeppert-Mayer is actually a few years older than that of Feynman. indeed yields (LABEL:Entropy_single_plus_loops). Note that the elements of the diagrams represent combinations of the measured cumulants - we do not have to know $J_{tt^{\prime}}$ or $H_{t}$ explicitly. We will sketch the basics of diagrammatics in appendix A.1 and refer to [18, 19, 20] for a more detailed description. Considering corrections to (LABEL:Approx_SingleBinEnt_Loops) by taking into account more diagrams can be beneficial in some cases, but we have found the resummed-loop approximation to be the most robust.
@@ III.2 Testing the approximation on artificial data
How does the dependence of $\mathrm{MIR}$ estimates on $\mathrm{dt}$ and $\Delta t$ in the regime of limited data looks like for other methods? In 4a, we demonstrate that assuming the data to be simply Gaussian, which amounts to setting $S=\ln\left(\det\left(c\right)\right)+\mathrm{const.}$ leads to a clear overestimation of the $\mathrm{MIR}$ (yellow bar). Also, just neglecting correlations between time bins is not feasible, as the green bar shows, indicating that actually about half of the information is captured by the interaction between bins. Characterizing this relation using the histogram method, assuming a quadratic dependence on $1/N_{R}$ and extrapolating to $1/N_{R}=0$, as suggested by Strong et al. [7], yields much better results (orange bar), but leads still to an overestimate. However, assuming a probability distribution of the form (LABEL:Def_generalized_Ising) for the probability distribution conditional on the stimulus and applying the approximation (LABEL:Entropy_single_plus_loops) (mixed Moba-MIRA) yields an excellent fit, provided that we remove the bias by subtracting the estimates obtained from shuffled data, see A.5 for details. Imposing the form of (LABEL:Def_generalized_Ising) for the output entropy as well (full Moba-MIRA) yields a slightly worse fit, however, still performs better than the method from [7].
Why does our approach, (LABEL:Entropy_single_plus_loops), work? In general pairwise probability distributions, (LABEL:Def_generalized_Ising), might not be appropriate because the data could be also shaped by higher-order interactions in addition. However, for the input entropy, the most important part of the covariance is due to the refractory period of the neurons, as visible in the auto-correlation of the spikes, cf. A1. In other words, given the stimulus, the activity in two time bins is correlated mostly because a spike in one time bin suppresses a spike in another one  [^3] Note also that this interaction is not directed in time - whenever there is a spike at some point in time, one knows that there could not have been a spike both before and after this time. This effective suppression of spikes in neighbored time bins is an intrinsically two-point like interaction - at least as long as the time bin is not considerably shorter than the refractory period. It therefore does not come as a surprise that this approach works well for the input entropy. For the output entropy, however, we cannot make a similar argument and indeed, the pairwise approach clearly works worse in this case, in particular for smaller time bins.
@@ IV Results: Application of Moba-MIRA on retinal stimulus response
The results we have presented so far were obtained with the “full” version of Moba-MIRA, that is, we have used the approximation (LABEL:Entropy_single_plus_loops) for the input and the output entropies. However, in the analysis of the artificial data, we have seen that mixed Moba-MIRA yields results visibly closer to the ground truth, cf. 4a. Also, we have found single ON-cells for which full Moba-MIRA seems to deliver unreliable results, see A.5. Luckily, the estimates for the $\mathrm{MIR}$ of our real data typically already converge at values for $\Delta t$ as low as $80\mathrm{ms}$, see A2. In this case $k=8$ so that it is feasible to compute the estimate for $\mathrm{MIR}$ also in the mixed-Moba-MIRA version. This allows us to compare our results shown in 5 and 6, obtained by full Moba-MIRA, by their counterparts obtained by mixed Moba-MIRA, which we show in III.2. We observe that there are no qualitative differences between these two approaches.
@@ V Discussion
Also, the pairwise model (LABEL:Def_generalized_Ising) provides even more potential to reduce numerical costs: By using the alternative definition
@@
for the mutual-information rate, one can reduce the numerical cost from being quadratic in $k$ to linear. Indeed, with the definition (LABEL:MIR_def_proper), one has to compute the correlations of $\frac{k\left(k-1\right)}{2}$ pairs of bins, assembled from the $k$ time bins in the time window $\Delta t$. With the definition (LABEL:MIR_diff_Def), however, one only has to compute the correlation of the activity in the most advanced time bin, corresponding to the interval $\left(\Delta t,\Delta t+\mathrm{dt}\right]$, with the $k$ bins inside $\left[0,\Delta t\right]$, therefore only $k$ correlations. As noted before [11], both definitions for the MIR are equivalent when the correlations between time bins decay to 0. In practice, the achievable $\Delta t$ might not always be large enough to make both expressions agree, but even then, (LABEL:MIR_diff_Def) is an appealing alternative because it is closer to a differential definition of a rate or flux.
@@
In this work we applied Moba-MIRA to estimate mutual-information rates of individual neurons. Our method can however be extended to account for populations by modeling the correlation between different neurons at different times, going beyond the study of zero-time-lag correlations and their impact on information [47, 48, 17, 49]. Even if undersampling might be an issue there, we expect Moba-MIRA to be very useful, as methods based on histogram approaches would require an even larger amount of data, often beyond existing experimental datasets. With Moba-MIRA, however, only the estimation of correlations matrices is required, which reduces the necessary dataset size. We thus expect that reliable estimates can be given at least for pairs of neurons, analyzed with a temporal resolution comparable to that employed in this study. Currently, extending Moba-MIRA in that direction can be hindered by the lack of ground truth estimation for large populations, and because of this we leave it for future developments.
@@ Appendix A Technical details and additional analysis
where by $\left\langle\dots\right\rangle_{P}$ we denote the average over the empirical distribution. The function $H_{i}$ is arbitrary in our technical framework - one might choose, e.g., $H_{i}\left(n\right)=-\ln\left(n!\right)-e^{h_{i}}$ in case one would like to expand around a Poissonian theory. We, instead, implement a maximum-entropy approach choosing it as power series according to (LABEL:Hamiltonian_infinitely_many_coeff). With this choice, we force our statistical model to reproduce all empirically measured cumulants for single bins. This is ensured by choosing the interaction matrix $J$ and the (infinitely many) parameters of $\boldsymbol{H}$, $\lambda_{1},\lambda_{2},\dots$ accordingly. Because $\ln\left({\cal Z}\right)$ is the cumulant-generating function, we can express this condition as